A rock and a half rock weigh a pound and a half. The rock and the half rock weigh the same. How much would half the whole rock and the whole of the half rock weigh?
a 2 1/2 b 1 1/8 c 1 7/2 d 2 1/4 e None of these.

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(30993)

A rock and a half rock weigh a pound and a half. The rock and the half rock weigh the same. How much would half the whole rock and the whole of the half rock weigh?
a 2 1/2 b 1 1/8 c 1 7/2 d 2 1/4 e None of these.
----------------
The rock and 1/2 rock weigh 3/4 pounds each.
Adding 1/2 a rock increased the total by 3/8 pounds
1.5 + 3/8 = 1 7/8 pounds
None of those, unless there's a typo in C

Answer by richwmiller(9143)

let x= whole rock
y/2=half rock
x+y/2=1.5=3/2
we are told that the whole rock and the half rock weight the same
x=y/2
x=3/2*1/2=3/4
y/2=3/4
y=6/4=3/2
half the whole rock =3/8
the whole of half rock =3/4
3/8+3/4=
3/8+6/8=
9/8= 1 1/8
answer b 1 1/8
we are not adding half of the whole rock but removing it.
we are removing 3/8
so instead of 3/4+3/4 we have 3/8 +3/4